# Exploring the Non-Markovian Dynamics in Depolarizing Maps

### v1

@@ IV.2 Non-Markovianity through visualizing the trajectory traced by map in the parameter space
The diagonalized form of the matrix elements of the intermediate map $\Phi(t,s)$ can be defined by three real parameters $\lambda_{1}(p),\lambda_{2}(p)$, and $\lambda_{3}(p)$, as shown in Eqs. (11) and (12). An arbitrary Pauli map is defined as [40, 23]
$\Phi(\rho)$ $=\frac{1}{2}\left(tr(\rho)\mathbb{I}+\sum_{i=1}^{3}\lambda_{i}(p)tr(\sigma_{i}\rho)\sigma_{i}\right)$
$=\begin{pmatrix}1&0&0&0\\ 0&\lambda_{1}(p)&0&0\\ 0&0&\lambda_{2}(p)&0\\ 0&0&0&\lambda_{3}(p)\end{pmatrix},$ (47)
where $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$ is a regular set of Pauli operators. Following [40, 23] the map $\Phi(\rho)$ is positive if $-1\leq\lambda_{i}\leq 1$ (cube in the parameter space) and the map $\Phi(\rho)$ is CP if $1\pm\lambda_{3}(p)\geq|\lambda_{1}(p)\pm\lambda_{2}(p)|$. The smooth process $\Phi(\rho)$ can then be determined by a continuous trajectory ${\bf{\lambda}}(p)$ in the parameter space. Such a trajectory provides a pictorial representation of the dynamical map in $\mathbb{R}^{3}$. By analysing the process trajectory in the parameter space, its (non-)Markovian properties are revealed. In [23], the dynamical maps were visualised via paths in the parameter space and their divisibility was investigated. The dynamics looses CP-divisibility under infinitesimal perturbations. Any qubit Pauli map $\Phi(\rho)$ between finite dimensional spaces can be described by three real parameters $\left(\lambda_{1}(p),\lambda_{2}(p),\lambda_{3}(p)\right)$ [19, 41, 42, 43]. Investigating the process trajectory in the parameter space brings out its divisibility properties. Our goal here is to detect the non-Markovianity of the depolarizing map in terms of its trajectory in the parameter space [13, 44, 23].
For any linear map, $\Phi(t,0)$, the trajectory $\{\lambda_{1},\lambda_{2},\lambda_{3}\}$ can be an arbitrary smooth curve inside the tetrahedron $1\pm\lambda_{3}\geq|\lambda_{1}\pm\lambda_{2}|$, Fig. (7). The curve starts initially, at $p=0$, from the corner $(\lambda_{1}(0)=1,\lambda_{2}(0)=1,\lambda_{3}(0)=1)$. If the traced trajectory curve stays in the tetrahedral part of the positive octant made from the cube shown in Fig. 7, the corresponding map $\phi_{t}$ is said to be CP. The trajectory of the curve residing in this positive octant pertains to some map $\phi_{t}$ with positive decoherence rates. The curve traversing out of this octant is indicative of a negative decoherence rate and non-CP behaviour [23].
@@
Figure 9: Tragectory-based visualization of dynamics (a) for $\alpha=0$; (b) for $\alpha=0.7$.
To discover direction during the dynamical progress in the parameter space, we use the following formula [23]
$\dot{\Phi}o\Phi^{-1}\left(\rho\right)=\frac{1}{2}\sum_{i=1}^{3}\frac{\dot{\lambda_{i}}(p)}{\lambda_{i}(p)}tr(\rho)\sigma_{i}$ (48)
@@
where, $\dot{\lambda}_{i}(p)\equiv\frac{d{\lambda}_{i}(p)}{dp}$. The vector $\bf{A}(p)$ can be established at any value $p$, highlighting the non-Markovian property of the dynamics. The dynamic is CP-divisible if and only if the vector $\bf{A}(p)$ drawn from the corner $(1,1,1)$ of the parameter space points inside the tetrahedron in Fig. (8). The scalar products of $\bf{A}(p)$ with vectors $(-1,1,1)$, $(1,-1,1)$, and $(1,1,-1)$ are all non-positive
@@
If a vector $\bf{A}(p)$ fails to satisfy Eq. ((50a)-(50c)), it indicates the non-Markovian CP-indivisible nature of the dynamics. For a general depolarizing map, its trajectory is given by $\lambda(p)=(\lambda_{1}(p),\lambda_{2}(p),\lambda_{3}(p))=\frac{3}{4}\alpha p^{2}-\alpha p-p+1$. In the case of Markovian dynamics, where $\alpha=0$ (no perturbation), the expressions become $A_{1}(p)=A_{2}(p)=A_{3}(p)=\left(1-p\right)$, and $\dot{\lambda}_{1}(p)=\dot{\lambda}_{2}(p)=\dot{\lambda}_{3}(p)=-1$. This results in $A_{1}(p)=A_{2}(p)=A_{3}(p)=\frac{-1}{1-p}<0$, for $p\in[0,1]$. The trajectory curve, depicted in Fig. (8a), is inside the positive octant; the curve traverses a path from $(1,1,1)$ to $(0,0,0)$. Consequently, the map $\Phi(t,0)$ satisfies Eq. (50a)-(50c), indicating that the dynamics is CP-divisible. Conversely, when $\alpha=0.7$ (with perturbation), the curve is modified to visualise the absolute value of eigenvalues. It is then observed that the curve traverses a reverse path for $p>\alpha_{-}(\alpha=0.7)$. This shift is shown in the plot given in Fig. (8b). The expressions for $A_{1}(p)$, $A_{2}(p)$, and $A_{3}(p)$ become $\left(\frac{42p-68}{21p^{2}-68p+40}\right)$. The inequalities in Eq. (42)-(48) are violated for the intermediate map when $0.8\leq p\leq 1$. Hence, $\Phi(t,0)$ is non-Markovian in this interval, in consistency with our findings in the previous section.
V Conclusion
@@
The notion of non-Markovianity in completely positive depolarizing noise is explored using various dynamical and geometrical techniques. The introduced non-Markovianity parameter ($\alpha$) of the depolarizing map, is the cause of perturbations and non-Markovianity. It contributes to breaking the CP divisibility of the intermediate map, assuring non-Markovianity. The non-Markovian behaviour of the depolarizing map could be identified by using appropriate witness operators. The intermediate evolution admits information backflow, showing non-monotonic variation in quantum memory, which could be justified using witness $X_{\chi(\alpha,q,p)}$. The witnessed non-Markovianity is quantified using two techniques: (a) BLP measure based on the distinguishability of states, (b) HCLA measure based on negativity of decoherence rates. We further study geometric visualisation techniques to gain a comprehensive understanding of non-Markovianity. It is observed that non-Markovianity contributes to enhancing the accessible volume of states during evolution. Furthermore, the trajectory traced during evolution in parameter space presents the loss of the CP divisibility under perturbations caused by non-zero $\alpha$.

### v2

@@ IV.2 Non-Markovianity through visualizing the trajectory traced by map in the parameter space
The diagonalized form of the matrix elements of the intermediate map $\Phi(p,q)$ can be defined by three real parameters $\lambda_{1}(p),\lambda_{2}(p)$, and $\lambda_{3}(p)$, as shown in Eqs. (11) and (12). An arbitrary Pauli map is defined as [68, 48]
$\Phi(\rho)$ $=\frac{1}{2}\left(\text{tr}(\rho)\mathbb{I}+\sum_{i=1}^{3}\lambda_{i}(p)\text{tr}(\sigma_{i}\rho)\sigma_{i}\right)$
$=\begin{pmatrix}1&0&0&0\\ 0&\lambda_{1}(p)&0&0\\ 0&0&\lambda_{2}(p)&0\\ 0&0&0&\lambda_{3}(p)\end{pmatrix}.$ (38)
Following [68, 48] the map $\Phi(\rho)$ is positive if $-1\leq\lambda_{i}\leq 1$ (cube in the parameter space) and the map $\Phi(\rho)$ is CP if $1\pm\lambda_{3}(p)\geq|\lambda_{1}(p)\pm\lambda_{2}(p)|$. The smooth process $\Phi(\rho)$ can then be determined by a continuous trajectory ${\bf{\lambda}}(p)$ in the parameter space. Such a trajectory provides a pictorial representation of the dynamical map in $\mathbb{R}^{3}$. By analyzing the process trajectory in the parameter space, its (non-)Markovian properties are revealed. In [48], the dynamical maps were visualized via paths in the parameter space and their divisibility was investigated. The dynamics loses CP divisibility under infinitesimal perturbations. Any qubit Pauli map $\Phi(\rho)$ between finite dimensional spaces can be described by three real parameters $\left(\lambda_{1}(p),\lambda_{2}(p),\lambda_{3}(p)\right)$ [20, 69, 70, 71]. Investigating the process trajectory in the parameter space brings out its divisibility properties. Our goal here is to detect the non-Markovianity of the depolarizing map in terms of its trajectory in the parameter space [13, 72, 48].
For any linear map, $\Phi(t,0)$, the trajectory $\{\lambda_{1},\lambda_{2},\lambda_{3}\}$ can be an arbitrary curve inside the tetrahedron $1\pm\lambda_{3}\geq|\lambda_{1}\pm\lambda_{2}|$; Fig. 8. The curve starts initially, at $p=0$, from the corner $(\lambda_{1}(0)=1,\lambda_{2}(0)=1,\lambda_{3}(0)=1)$. If the traced trajectory curve stays in the tetrahedral part of the positive octant made from the cube shown in Fig. 7, the corresponding map $\phi_{t}$ is said to be CP. The trajectory of the curve residing in this positive octant pertains to some map $\phi_{t}$ with positive decoherence rates. The curve traversing out of this octant is indicative of a negative decoherence rate and non-CP behavior [48].
@@
Figure 9: Trajectory-based visualization of dynamics (a) for $\alpha=0$ and (b) for $\alpha=0.7$.
To discover direction during the dynamical progress in the parameter space, we use the following formula [48]:
$\dot{\Phi}\circ\Phi^{-1}\left(\rho\right)=\frac{1}{2}\sum_{i=1}^{3}\frac{\dot{\lambda_{i}}(p)}{\lambda_{i}(p)}tr(\rho)\sigma_{i}$ (39)
@@
where $\dot{\lambda}_{i}(p)\equiv\frac{d{\lambda}_{i}(p)}{dp}$. The vector $\bf{A}(p)$ can be established at any value $p$, highlighting the non-Markovian property of the dynamics. The dynamics is CP divisible if and only if the vector $\bf{A}(p)$ drawn from the corner $(1,1,1)$ of the parameter space points inside the tetrahedron in Fig. 8. The scalar products of $\bf{A}(p)$ with vectors $(-1,1,1)$, $(1,-1,1)$, and $(1,1,-1)$ are all nonpositive,
@@
If a vector $\bf{A}(p)$ fails to satisfy Eq. (41a)-(41c), it indicates the non-Markovian CP-indivisible nature of the dynamics. For a general depolarizing map, its trajectory is given by $\lambda(p)=(\lambda_{1}(p),\lambda_{2}(p),\lambda_{3}(p))=\frac{3}{4}\alpha p^{2}-\alpha p-p+1$. In the case of Markovian dynamics, where $\alpha=0$ (no perturbation), the expressions become $A_{1}(p)=A_{2}(p)=A_{3}(p)=\left(1-p\right)$, and $\dot{\lambda}_{1}(p)=\dot{\lambda}_{2}(p)=\dot{\lambda}_{3}(p)=-1$. This results in $A_{1}(p)=A_{2}(p)=A_{3}(p)=\frac{-1}{1-p}<0$, for $p\in[0,1]$. The trajectory curve, depicted in Fig. (8a), is inside the positive octant; the curve traverses a path from $(1,1,1)$ to $(0,0,0)$. Consequently, the map $\Phi(p,0)$ satisfies Eqs. (41a)-(41c), indicating that the dynamics is CP divisible. Conversely, when $\alpha=0.7$ (with perturbation), the curve is modified to visualize the absolute value of the eigenvalues. It is then observed that the curve traverses a reverse path for $p>\alpha_{-}(\alpha=0.7)$. This shift is shown in the plot given in Fig. 8(b). The expressions for $A_{1}(p)$, $A_{2}(p)$, and $A_{3}(p)$ become $\left(\frac{42p-68}{21p^{2}-68p+40}\right)$. The inequalities in Eqs. (37)-(39) are violated for the intermediate map when $0.8\leq p\leq 1$. Hence, $\Phi(p,0)$ is non-Markovian in this interval, consistent with our findings in the previous section.
V Brief overview of depolarizing noise in higher dimensional systems
An understanding of non-Markovianity was attempted in previous sections using several dynamical and geometric characterization techniques at the qubit level. Here, a foray is made into non-Markovianity exhibited by higher dimensional systems, evolving through the depolarizing channel, using a dynamical and geometrical characterization method.
We define non-Markovian depolarizing noise in an $N$-level qudit system [73] using the following Kraus operator
$E_{r,s}=\begin{cases}\sqrt{\left(1-\frac{(N^{2}-1)}{N^{2}}\alpha p\right)\left(1-\frac{(N^{2}-1)}{N^{2}}p\right)}\mathbb{I}_{N}&\text{for $r=0\And s=0$};\\ \sqrt{1+\alpha\left(1-\frac{(N^{2}-1)}{N^{2}}p\right)\frac{p}{N^{2}}}U_{r,s}&\text{for $0\leq r,s\leq(N-1)$, $(r,s)\neq(0,0)$ },\end{cases}$ (42)
where $U_{r,s}$ is the Weyl operator defined as
$U_{r,s}=\sum_{i=0}^{N-1}\omega^{ir}_{N}\ket{i}\bra{i\oplus s},0\leq r,s\leq(N-1),$ (43)
where $\omega_{N}=\exp{\left(\frac{2\pi\iota}{N}\right)}$, and $\oplus$ denotes addition modulo $N$. These Kraus operators are now used in the study of non-Markovianity in higher dimensional space. The N-level depolarizing map provides a complex example with singularities, and non-Markovianity. Understanding the variation in non-Markovianity and singularity structure with increased system dimension is a worthwhile exploration. An attempt is made in this direction for a three-level qutrit system, where the number of Kraus operators given in Eq. (42) is nine. Extensions to higher dimensions can be made in a similar fashion.
V.1 non-Markovianity for qutrit system (N=3)
We provide a perspective to non-Markovian effects in a qutrit system undergoing depolarizing noise. We present the dynamical characterization of these effects using the HCLA measure, which relies on negative decoherence rates. In addition the geometrical visualization of non-Markovian effects is also attempted using volume change in the physical states of the system.
V.1.1 HCLA Measure
For an $N$-level qudit system, $k(p)=p+\alpha p-\frac{(N^{2}-1)}{N^{2}}\alpha p^{2}$, [see Eq. (6)]. In the case of qutrit, it becomes $k(p)=p+\alpha p-\frac{8}{9}\alpha p^{2}$. Following the normalization procedure described in Sec. III.2, we obtain normalized rates. The integration of normalized rates from the point $\alpha_{-}$, obtained from solving the quadratic equation $09p+9\alpha p-8\alpha p^{2}-9=$ 0, to $1$ results in
[…]
=$=\Biggl{\{}\ln\left(p\right)+\ln\left(9+9\alpha-8p\alpha\right)\Biggr{\}}^{p=1}_{p=\alpha_{-}}.$
Figure 10 depicts the variation in non-Markovianity as a function of parameter $\alpha$. The increase in $\alpha$ results in higher non-Markovianity. Comparing the non-Markovianity in qubit and qutrit systems, one can observe a decrease with increase in dimension. The point $\alpha_{-}$ shifts to a higher value with increasing levels of the system, shrinking the region of negative decay rates. This observation further points to the vanishing of the characteristic singularity structure for systems with $N\geq 5$ for $p,\alpha$ ranging from 00 to $1$.
Figure 10: A plot of HCLA measure for qutrit system.
Figure 11: The variation in $||F_{3}||_{1}$ for $\alpha=0.7$.
V.1.2 Variation in volume of physical states
The geometric depiction of non-Markovian effects can be visualized using the variation in the volume of physical states, as in the previous Sec. IV.1. We extend this idea to the $N$-level system and also present the analysis of qutrit system (N=3). The generalized Gell-Mann matrix [45] satisfies the properties
$G_{0}=\frac{1}{\sqrt{d}}\mathbb{I}_{d};\quad G_{m}=G_{m}^{\dagger};\quad\text{tr}[G_{m}G_{n}]=2\delta_{mn},$ (45)
and can be used. For N=3, this results in nine basis operators given as
$G=\{\frac{1}{\sqrt{3}}\mathbb{I}_{3},\frac{1}{\sqrt{2}}G_{m}\},$ (46)
where $G_{m=1,2,..8}$ represents Gell-Mann matrices. Using these basis operators and the depolarizing channel map constructed using the operators defined in Eq. (42), we construct the $F_{3}$ matrix [74, 75], whose elements are
$F_{kl}:=\frac{1}{3^{2}}\text{tr}[G_{k}\mathcal{E}_{p}(G_{l})].$ (47)
The non-monotonic variation in $||F_{3}||_{1}$ is characteristic of non-Markovian dynamics and is depicted in Fig. 11 for the qutrit system, by the nonmonotonic behavior. Similar results can be obtained for higher dimensional systems.
((a))
((b))
Figure 12: The variation in $||\chi(\alpha,q,p)||_{1}$ with $\alpha=0.9,q=0.4$, for a (a) multiqubit map, and (b) multi-level map, highlights the impact on the singularity structure, which is invariant for a multiqubit map and changes for a multi-level map affecting non-Markovianity.
VI Investigating the non-Markovianity of depolarizing noise in multi-qubit Systems
In the previous section, we extended our investigation to encompass higher-dimensional systems, examining their non-Markovian behavior as they evolve through the depolarizing channel using both dynamical and geometric characterization approaches. We now turn our attention to the non-Markovianity observed in multiqubit systems. We examine the scenario in which qubits are exposed to independent depolarizing channels that can be adjusted to display both Markovian and non-Markovian dynamics [76]. Then, the dynamical evolution of the multi-qubit systems will be obtained by utilizing the Kraus operators constructed from the tensor product of single qubit Kraus operators [Eq. II]. Following the previously used method in Sec. III.1 and Appendix A, one can construct the Choi matrices $\chi(\alpha,q,p)$ of multiqubit maps.
The variation in trace norm of $\chi(\alpha,q,p)$ is useful to probe the singularity structure and NCP nature of higher-qubit and higher-level maps. The numerical variation in the trace norm of $\chi(\alpha,q,p)$ is shown in Fig. 12(a), depicting the invariant singularity structure due to the same $\alpha_{-}$ in the case of one-, two-, and three-qubit maps. Similarly, Fig. 12(b) reveals the increase in value of $\alpha_{-}$, indicating the possibility of the vanishing singularity structure and non-Markovianity with increasing levels of quantum system. Further, from the analysis of the two-qubit channel, using the right derivative of the trace norm of $\chi(\alpha,q,q+\epsilon)$ results in the function [7]
[…]
This reveals enhanced non-Markovian features in the two-qubit map compared to the single-qubit map. The variation of $g(q,\alpha)$ is depicted in Fig. 13. The two-qubit map reaches higher values of the function $g(q,\alpha)$ compared to the single-qubit map, indicating enhanced non-Markovianity.
Figure 13: The variation in $g(q,\alpha)$ for $\alpha=0.9$.
VII Conclusion
@@
The notion of non-Markovianity in completely positive depolarizing noise is explored using various dynamical and geometrical techniques. The introduced non-Markovianity parameter ($\alpha$) of the depolarizing map, is the cause of perturbations and non-Markovianity. It contributes to breaking the CP divisibility of the intermediate map, assuring non-Markovianity. The non-Markovian behavior of the depolarizing map could be identified by using appropriate witness operators. The intermediate evolution admits information backflow, showing non-monotonic variation in quantum memory, which could be justified using witness $X_{\chi(\alpha,q,p)}$. The witnessed non-Markovianity is quantified using two techniques: (a) BLP measure based on the distinguishability of states, and (b) HCLA measure based on the negativity of the decoherence rates. We further study geometric visualization techniques to gain a comprehensive understanding of non-Markovianity. It is observed that non-Markovianity contributes to enhancing the accessible volume of states during evolution. Furthermore, the trajectory traced during evolution in parameter space presents the loss of the CP divisibility under perturbations caused by non-zero $\alpha$. The extension of this analysis to higher dimensional systems is made and illustrated for a qutrit system. The non-Markovianity and singularity structure are observed to decreases with increase in system size. Finally, this analysis is extended to multi-qubit systems, where it is observed that non-Markovianity increases with the number of qubits. The singularity structure remains invariant in multiqubit maps; however, it is affected and ultimately lost in higher-dimensional (qudit) maps.
Appendix A Constructing the Choi matrix $\chi(\alpha,q,p)$ for the intermediate map
In this appendix, we present the detailed mathematical construction of the Choi matrix for the intermediate map (propagator). We begin by constructing $U_{2\leftrightarrows 3}[\Phi(p,q)\otimes\mathbb{I}_{N^{2}}]U_{2\leftrightarrows 3}$ where $U_{2\leftrightarrows 3}$, for $N$-dimensional system, is the commutation (or “swap”) matrix [58, 56] between the “second” and the “third” subspaces [7]. Second, we apply $U_{2\leftrightarrows 3}[\Phi(p,q)\otimes\mathbb{I}_{N^{2}}]U_{2\leftrightarrows 3}$ on $\rm{vec}(\ket{\Psi}\bra{\Psi})$, where |$\ket{\Psi}=\sum_{i=0}^{N-1}\frac{1}{\sqrt{N}}\ket{ii}$. $U_{2\leftrightarrows 3}$ is written as $\mathbb{I}_{N}\otimes U_{P}\otimes\mathbb{I}_{N}$, where $U_{P}$ is a permutation matrix satisfying the property $U_{P}(A\otimes B)U_{P}=B\otimes A$, and is defined as
$U_{P}=\sum_{k=0}^{N-1}\sum_{l=0}^{N-1}\ket{k}\bra{l}\otimes\ket{l}\bra{k}.$ | . (49)
We assume that $\phi(p,q)\otimes\mathbb{I}_{N^{2}}$ acts as a tensor product of four spaces with the same dimension, $\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\mathcal{H}_{3}\otimes\mathcal{H}_{4}$. Then $U_{2\leftrightarrows 3}$ denotes the permutation matrix interchanging the second and third subspace, $i.e.$, $U_{2\leftrightarrows 3}=\mathbb{I}_{N}\otimes U_{P}\otimes\mathbb{I}_{N}$ Lastly, we write the result as a matrix, $i.e.$, “devectorize” to construct the Choi matrix of the intermediate map as follows
$\chi(\alpha,q,p)=\left[\Phi(p,q)\otimes\mathbb{I}\right]\ket{\Psi}\bra{\Psi}.$ | . (50)
Thus, we obtain the matrix in Eq. (11) for the case of $N=2$.